OpenStudy (anonymous):

if I need a pink shade of paint which consists of red and white paint. i need 128 gallons of this pink paint. the amount of red is three times the amount of white. how many gallons of each paint do i need?

OpenStudy (anonymous):

Let the amount of white paint be = x then, amount of red paint is = 3x so. 3x + x = 128 4x = 128 x = 128/4 = 32 amount of red paint needed = 3*32 = 96 gallons amount of white paint needed = 32 gallons

OpenStudy (matthewrlee):

Go to Home Depot. They can help.